# Beyond Maximum Likelihood: Variational Inequality Estimation for Generalized Linear Models

### v1

@@ 1.1 Notation
2 Background and Problem Setup
We begin by reviewing the GLM framework and establishing notation used throughout the paper. We then introduce the MLE approach and its VI reformulation, which serves as the foundation for our theoretical analysis.
Given i.i.d. samples $(\bm{x},y)$, where $\bm{x}\in\mathbb{R}^{d}$ denotes the vector of regressors and $y\in\mathbb{R}$ denotes the response, GLM specifies the conditional distribution of response $y\mid\bm{x}$ using the exponential family. The mean response is related to a linear predictor through a specified link function $g:\mathbb{R}\to\mathbb{R}$:
where $\bm{\beta}=(\beta_{0},\dots,\beta_{d})\in\mathbb{R}^{d+1}$ is the coefficient vector with the intercept term. GLMs include many classical models as special cases. For instance, logistic regression corresponds to a GLM with a binomial distribution and a logit link, while Poisson regression corresponds to a GLM with a Poisson distribution and a logarithm link function.
A standard approach for estimating $\bm{\beta}$ is the MLE, which maximizes the log-likelihood of the observed data. Given a statistical model with probability density (or mass) function $p(y;\bm{\beta})$ and a dataset $\{y^{i}\}_{i=1}^{N}$, the MLE is defined as
$\max_{\bm{\beta}\in\mathbb{R}^{d+1}}\;\prod_{i=1}^{N}p(y^{i};\bm{\beta}),$
which is equivalently obtained by minimizing the empirical negative log-likelihood (NLL):
$\min_{\bm{\beta}\in\mathbb{R}^{d+1}}\;\sum_{i=1}^{N}-\log p(y^{i};\bm{\beta}).$
In the GLM framework, this objective can be expressed in the general empirical risk form with loss function $\ell$ and link function $g$:
$\min_{\bm{\beta}\in\mathbb{R}^{d+1}}\;\mathcal{L}_{N}(\bm{\beta}):=\sum_{i=1}^{N}\ell\left(g^{-1}\left(\beta_{0}+\sum_{j=1}^{d}\beta_{j}x^{i}_{j}\right),y^{i}\right).$ (MLE)
A variety of numerical algorithms can be employed to solve (MLE). Classical second-order methods, including Newton-type schemes such as Fisher scoring and iteratively reweighted least squares (IRLS) [green1984iteratively, mccullagh1989generalized], are commonly used in GLMs, while large-scale problems are often addressed using first-order approaches such as gradient descent or its stochastic variants [nesterov2013introductory, kushner2003stochastic].
In classical settings such as linear, logistic, and Poisson regression with canonical link functions, the optimization problem (MLE) is convex and smooth, ensuring global optimality and fast convergence of standard algorithms. However, when the GLM employs a non-canonical link, the optimization landscape can become ill-posed, leading to slow convergence of these numerical methods; see, e.g., [he2020point]. In addition, if the inverse link function is non-monotone or non-smooth (see, e.g., [cao2015poisson], where in practice one has to impose a lower bound on the Poisson intensity function), the resulting NLL objective is generally non-convex and non-smooth. Such situations arise in practice, for example, in Poisson regression with a clipped exponential link or in non-canonical extensions of GLMs that use piecewise linear activations such as hinge or ReLU functions. To address these challenges, it is natural to consider reformulating the estimation problem from an operator perspective, one that captures the equilibrium condition defining the MLE but relaxes convexity and differentiability requirements.
To overcome these difficulties, we reformulate the estimation problem as a variational inequality. The VI framework introduces an operator-based characterization that allows analysis and computation under weaker regularity assumptions. Let $V:\mathbb{R}^{d+1}\to\mathbb{R}^{d+1}$ denote the population operator
where we introduce the augmented feature vector $\tilde{\bm{x}}:=[1;\bm{x}]\in\mathbb{R}^{d+1}$. The target parameter $\bm{\beta}^{\star}$ is defined as the solution to the variational inequality
$\langle V(\bm{\beta}^{\star}),\,\bm{\beta}-\bm{\beta}^{\star}\rangle\geq 0,\quad\forall\,\bm{\beta}\in\mathbb{R}^{d+1}.$ (2.2)
Given i.i.d. observations, $\{(\bm{x}^{i},y^{i})\}_{i=1}^{N}$, the empirical VI estimator $\hat{\bm{\beta}}_{N}$ is obtained by the empirical average:
where
$V_{N}(\bm{\beta}):=\frac{1}{N}\sum_{i=1}^{N}\!\left(g^{-1}\left(\beta_{0}+\sum_{j=1}^{d}\beta_{j}x^{i}_{j}\right)-y^{i}\right)\tilde{\bm{x}}^{i}.$ (2.3)
The VI estimator generalizes the first-order optimality condition of the MLE and reduces to it under canonical link functions, implying that standard GLM estimation can be viewed as an implicit VI problem. This formulation was first considered by Juditsky and Nemirovski [juditsky2019signal] in the context of nonlinear least-squares estimation for GLMs. Their approach builds upon the seminal ideas of Rosenblatt’s perceptron algorithm [rosenblatt1958perceptron] and subsequent perceptron-like methods [kalai2009isotron, kakade2011efficient]. Since then, the VI framework has been extended to a variety of settings, demonstrating its versatility and effectiveness [juditsky2020convex, juditsky2023generalized, cheng2024point, zhou2024nonlinear, lou2025accurate]. Compared with the nonlinear least-squares formulation, which often leads to non-convex objectives under nonlinear link functions, the VI approach assumes only that the link function is strongly monotone, thereby inducing a strongly convex optimization structure and enabling stable computation. Building upon these foundational developments, our work provides a unified theoretical and algorithmic treatment of VI-based estimation for GLMs, establishing its statistical optimality and computational convergence properties.
2.1 Summary of main results
We develop a rigorous theoretical and algorithmic analysis of the proposed VI-based estimator for GLMs, establishing both statistical guarantees and computational convergence results. In contrast to existing results such as [juditsky2019signal], which primarily focus on using stochastic approximation schemes based on VI to solve nonlinear least-squares problems, our work highlights the statistical advantages and theoretical robustness of a VI-based estimator relative to the MLE for GLMs. The main contributions are summarized as follows:
  • Statistical guarantees. Beyond the non-asymptotic convergence results established for stochastic approximation schemes, we derive both finite-sample estimation error bounds and asymptotic normality for the sample-average approximation solution of the VI, extending the stochastic VI framework of [king1993asymptotic, xu2010sample]. Specifically, the estimator achieves the optimal finite-sample rate and asymptotic normality:
    $\|\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star}\|\leq\tilde{\mathcal{O}}\left(\sqrt{\frac{d}{N}}\right)\ \mbox{w.h.p},\quad\text{and}\quad\sqrt{N}(\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star})\ \xrightarrow{d}\ \mathcal{N}(\mathbf{0},\Sigma^{\star}),$
    for some covariance matrix $\Sigma^{\star}$ that depends on the true parameter and the data distribution. We also establish linear convergence of fixed-point iterative algorithms for solving the empirical VI and the standard sublinear convergence rate for stochastic approximation under streaming data.
  • Extension to general link functions. We extend the theoretical results within the stochastic VI framework to accommodate non-smooth, potentially non-monotone link functions that satisfy the strong Minty condition, such as ReLU and clipped links. These link functions can be used to handle, for instance, heavy-tailed data, stabilize variance, and improve robustness to outliers, thereby broadening the applicability of the proposed framework beyond the assumptions of existing methods. In contrast, achieving comparable performance with MLE requires the gradient of the composite mapping—i.e., the loss function composed with the inverse link—to satisfy a gradient-based strong Minty condition, which involves an additional application of the chain rule. For models with non-canonical link functions, this condition is often difficult to satisfy, resulting in a flatter optimization landscape and slower convergence. When non-smooth or non-convex structures are involved, MLE algorithms may further become trapped in suboptimal solutions and suffer from numerical instability due to the failure of the chain rule.
3 From MLE to VI
We consider the GLM defined in (2.1), where the response variable follows an exponential-family distribution:
$y\mid\bm{x}\sim\mathrm{Exponential\,Family}(\mu),\quad\mu=g^{-1}(z),\quad z=\tilde{\bm{x}}^{\top}\bm{\beta}.$
In the formulation (MLE), the specification of (i) the loss function $\ell$ and (ii) the inverse link function $g^{-1}$ arises from modeling assumptions about the data-generating process.
3.1 Equivalence under canonical links
When the model employs a canonical link function, the gradient of the NLL function in (MLE) coincides exactly with the operator defining the variational inequality in (VI), i.e.,
$\nabla\mathcal{L}_{N}(\bm{\beta})=V_{N}(\bm{\beta}).$ (3.1)
This equivalence has been noted in [juditsky2019signal] for logistic regression. For instance, consider binary logistic regression, where the empirical NLL is given by
$\mathcal{L}_{N}(\bm{\beta})=-\frac{1}{N}\sum_{i=1}^{N}\left(y^{i}\log g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})+(1-y^{i})\log(1-g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i}))\right),$
@@ 3.1 Equivalence under canonical links
$\nabla\mathcal{L}_{N}(\bm{\beta})=\frac{1}{N}\sum_{i=1}^{N}\frac{(g^{-1})^{\prime}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})}{g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})\,(1-g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i}))}\left(g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})-y^{i}\right)\tilde{\bm{x}}^{i}.$
@@
Such an equivalence also holds for linear, exponential, and Poisson regression models, indicating that under canonical link functions, the MLE and VI formulations share identical first-order optimality conditions. In contrast, when a non-canonical link function is used, the gradient of the likelihood and the VI operator differ, leading to distinct estimation dynamics and potentially different convergence behaviors. The similar identity holds for other types of link functions such as linear, Poisson, and exponential; representative examples are summarized in Table 1 and the derivations can be found in Appendix A.
Table 1: MLE and VI Formulations under Different Links
Model Loss Function $\ell(u,y)$ Inverse Link $g^{-1}(z)$ $\nabla\mathcal{L}(\bm{\beta})=V(\bm{\beta})?$
Logistic $-y\log u-(1-y)\log(1-u)$ $e^{z}/(1+e^{z})$
$1/2+\arctan(z)/\pi$
Normal $\frac{1}{2}(u-y)^{2}$ $z$
$\max\{0,z\}$
Exponential $\log u+y/u$ $z^{-1}$
$e^{-z}$
Poisson $-y\log u+u$ $e^{z}$
$\log(1+e^{z})$
$\max\{c,\min\{e^{z},C\}\}$
In the following non-canonical settings, the MLE landscape can become very flat, leading to slow convergence. Moreover, the loss function may be non-smooth or non-convex, violating standard subdifferential calculus rules [li2020understanding] and introducing additional challenges for both optimization and the statistical consistency of the MLE. Such cases arise, for example, in Poisson regression models with softplus or clipped exponential link functions.
3.2 Flatness of MLE optimization landscape
It is well known that MLE optimization algorithms can converge slowly near the true parameter, especially when the likelihood surface is nearly flat or ill-conditioned. We show that even when standard smooth link functions are applied to improve curvature, the resulting optimization landscape of the MLE can still exhibit considerably slower convergence than the VI formulation. This phenomenon is illustrated using the softplus link function, under which the MLE score and the VI operator define different estimating equations.
@@ 3.2 Flatness of MLE optimization landscape
$g^{-1}(z)=\log(1+e^{z}),\quad(g^{-1})^{\prime}(z)=\sigma(z):=\frac{1}{1+e^{-z}},\quad(g^{-1})^{\prime\prime}(z)=\sigma(z)\left(1-\sigma(z)\right).$

### v2

@@ Appendix A Notation
Appendix B Detailed MLE and VI Optimization Landscape Derivations
In this appendix, we supplement the motivation provided in Section 3 with detailed mathematical derivations and a rigorous analysis of the optimization landscapes. First, we explicitly derive the gradients for several classical GLMs, including logistic, linear, exponential, and Poisson regression, to demonstrate the precise equivalence between MLE and VI formulation under canonical link functions (Appendix B.1). Second, we investigate the pathological behaviors of MLE in non-canonical settings. We provide an analytical calculation of the Hessian for the softplus link to quantify the ill-conditioning (flatness) of the MLE objective, and we present a detailed characterization of the non-convexity and non-smoothness induced by robust clipped links. These analyses highlight the structural advantages of the VI operator, which retains monotonicity even when the corresponding likelihood objective fails to be convex.
B.1 Examples of Vector Fields for Canonical MLE and VI
B.1.1 Logistic regression
For the binary logistic regression, the empirical NLL is given by
$\mathcal{L}_{N}(\bm{\beta})=-\frac{1}{N}\sum_{i=1}^{N}\Bigl($ $y^{i}\log g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})+(1-y^{i})\log\bigl(1-g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})\bigr)\Bigr).$
@@ B.1.1 Logistic regression
$\nabla\mathcal{L}_{N}(\bm{\beta})=\frac{1}{N}\sum_{i=1}^{N}\;$ $\frac{(g^{-1})^{\prime}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})}{g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})\bigl(1-g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})\bigr)}\cdot\Bigl(g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})-y^{i}\Bigr)\,\tilde{\bm{x}}^{i}.$
@@
B.1.2 Linear regression
whose gradient is
B.1.3 Exponential regression
and the corresponding gradient is
B.1.4 Poisson regression
For the Poisson model, the empirical NLL is
The gradient is
B.2 Flatness of MLE optimization landscape
We show that even when standard smooth link functions are applied to improve curvature, the resulting optimization landscape of the MLE can still exhibit considerably slower convergence than the VI formulation. This phenomenon is illustrated using the softplus link function, under which the MLE score and the VI operator define different estimating equations.
@@ B.2 Flatness of MLE optimization landscape
$g^{-1}(z)=\log(1+e^{z}),\quad(g^{-1})^{\prime}(z)=\sigma(z):=\frac{1}{1+e^{-z}},\quad$
and
$(g^{-1})^{\prime\prime}(z)=\sigma(z)\left(1-\sigma(z)\right).$